# InfNeRF: Towards Infinite Scale NeRF Rendering with O(log n) Space Complexity

### v1

@@ Abstract
In this work, we extend this proven LoD technique to Neural Radiance Fields (NeRF) by introducing an octree structure to represent the scenes in different scales. This innovative approach provides a mathematically simple and elegant representation with a rendering space complexity of , aligned with the efficiency of mesh-based LoD techniques. We also present a novel training strategy that maintains a complexity of . This strategy allows for parallel training with minimal overhead, ensuring the scalability and efficiency of our proposed method. Our contribution is not only in extending the capabilities of existing techniques but also in establishing a foundation for scalable and efficient large-scale scene representation using NeRF and octree structures.
Keywords:
novel view synthesis radiance fields level of detail
1 Introduction
Recent progress in Neural Radiance Fields (NeRF) has significantly advanced the field of photo-realistic novel-view synthesis [25, 2, 4, 34]. Surpassing traditional 3D representations like meshes, NeRF demonstrates a superior capability in generating accurate novel views, particularly for transparent or reflective geometries. However, despite these advancements, a predominant focus of NeRF research has been confined to single objects or limited-scale scenes. The challenge of effectively representing extensive, large-scale scenes remains largely underexplored and unresolved.
@@ 1 Introduction
1.1 Block-NeRF and Mega-NeRF
@@
Towards this direction, recent large-scale NeRF approaches, such as Block-NeRF [32] and Mega-NeRF [36] divide the whole neural field equally in space into smaller, more manageable blocks, referred to as grid partition in this paper. This allows the rendering device to load only the data relevant to the camera’s current location. However, these methods encounter challenges in scenarios requiring a bird’s eye view of the entire scene. Such a panoramic view plays a crucial role in human interactive navigation. For example, to adjust the position of the camera, it is much more efficient to first zoom out, identify the target location, and then zoom in again, compared to the laborious process of navigating at a constant zoomed-in level. However, in such scenarios, the existing methods need to load all blocks simultaneously, which can overwhelm the memory capacity of the device. In addition to the memory issue, these approaches are prone to aliasing artifacts that arise from an insufficient sampling rate for high-frequency signals as illustrated in Fig.1.
1.2 Proposed InfNeRF
@@
In this paper, we propose InfNeRF, a novel method that addresses the above challenges. The core concept of InfNeRF is inspired by the classical mesh-based Level of Detail (LoD) techniques [12, 16, 29, 34]. An overview is illustrated in Fig.2. Specifically, we construct an LoD tree for the target scene, where each node encapsulates a small NeRF with a roughly equivalent number of parameters. The root node provides a coarse representation of the entire scene, while its successive children progressively describe more detailed sub-regions. During rendering, each sampling point is forwarded by a different NeRF, depending on its location and radius. The density and color then are accumulated to determine the pixel color.
@@
As illustrated in Fig. 3, InfNeRF only requires a minimum subset of the nodes in rendering, chosen based on their proximity to the camera and their intersection within the frustum. This significantly reduces the memory burden and the I/O time for retrieving parameters from disk or cloud storage.
The memory complexity is notably efficient at , where denotes the total information captured in the reconstruction. In the real world, with , we can represent the Earth down to a resolution of centimeters.
@@
One remarkable feature of InfNeRF is that the parent nodes in the LoD tree act as a smoother, low-pass-filtered version of the scene. Therefore, the proposed method naturally removes high-frequency components in distant views, which effectively reduces the aliasing artifacts and enhances the visual quality of the generated images (Fig. 1). In our experiments, InfNeRF achieves superior rendering quality for large-scene reconstruction, with an improvement of over 2.4dB in PSNR while accessing only 17% of the total parameters.
@@
In addition, to facilitate efficient training of InfNeRF, we introduce a novel training strategy that maintains a time and space complexity of . Moreover, this strategy allows for the training process to be effectively parallelized across machines.
InfNeRF distinguishes itself from existing methods such as Block-NeRF [32] and Mega-NeRF [36] in that it divides the scene in both spatial and scale dimensions while existing approaches typically only consider the spatial aspect. InfNeRF imposes no constraints on the underlying NeRF model and makes no assumptions about the scene, making it an adaptable framework that can seamlessly accommodate and leverage the rapid advancements in NeRF models.
@@
2 Related Work
2.0.1 Photogrammetry.
@@ 2 Related Work
Among the established reconstruction techniques, mesh-based reconstruction methods called photogrammetry have been used widely for it efficiency and robustness. This workflow includes successive steps: Structure from Motion (SfM) [1, 42, 30, 26] for camera pose estimation, Multi-View Stereo (MVS) techniques [18, 5, 31] for depth computation, mesh reconstruction [23, 39], and texture mapping [40]. These steps form the basic workflow to create a textured mesh that can be efficiently rendered by a traditional graphics pipeline. Open-source tools, such as COLMAP [30] and OpenMVS[7], have seamlessly integrated this workflow and gained widespread adoption in the academic community. Furthermore, commercial software solutions like ContextCapture and DJI Terra [14] have significantly elevated the robustness and efficiency to meet industrial standards.
2.0.2 Level of Detail.
@@
When the scene is large, it becomes impractical to store it entirely in memory. In this case, the Level of Detail (LoD) or Hierarchical Level of Detail (HLoD) [12, 16, 19, 20, 38, 24] provides a framework for efficiently storing, fetching, and rendering the scene.
@@
Take tiled 2D map [35] used widely in 2D navigation system for an example. The maps in different resolutions are cut into 256256 PNG images referred to as tiles. These tiles are organized into a quadtree structure, typically comprising 22 levels. This hierarchical arrangement allows web browsers to rapidly fetch data and render any location at any zoom level with constant space complexity.
@@
A similar technique can also be applied to 3d mesh. After the reconstruction of the scene into one detailed mesh, this textured mesh undergoes multiple downsampling [13, 41] resulting in a mesh pyramid with varying levels of detail, ranging from fine to coarse. Subsequently, these meshes are partitioned into numerous tiles and organized hierarchically, resembling a tree structure similar to a quadtree in 2D mapping. During rendering, open source 3D engines such as CesiumJS [8] or OpenSceneGraph [6] walk the tree, then progressively fetch tiles based on their proximity and intersection with the frustum. This LoD technique enables users to have a seamless 3D navigation experience in a very large scene while maintaining a low memory footprint.
2.0.3 NeRF and Large-Scale NeRF.
@@
NeRF [25] employs Multi-Layer Perceptrons (MLPs) for NeRF to capture a continuous density and viewpoint-dependent radiance of a scene. Plenoxel introduces a voxel-dense grid [17], which significantly improves the training and rendering speeds of NeRF, albeit with a significant VRAM requirement. Methods like TensoRF [10] and MERF [28] decompose the dense grid into low dimension feature tensor, while InstantNGP [27] opts to store these vectors in a sparse hash table. These strategies greatly enhance model compactness while maintaining speed.
2.0.4 Large-Scale Scene and Partition at Space Dimension.
@@
While previous NeRF approaches primarily focus on scenes of limited scale, scaling up NeRF to handle large-scale scenes, such as cities or even the entire Earth, would empower a broader range of applications. Mip-NeRF 360 [3] proposed a scene contraction technique to compress the unbounded scene into a bounded region. However, this approach compromises quality in the compressed areas. Grid-Guided NeRF [45] extends its capacity by combining a 2D feature grid and MLP, but its reliance on a 2D plane assumption limits its capability to handling complicated 2D scene. Block-NeRF [32], Mega-NeRF [36], SMERF [15] and ScaNERF [43] divide the scene into several sub-regions and reconstruct them separately. In its specialized indoor scenario, SMERF can even achieve impressive real-time rendering. However, all these methods suffer from a common issue: when rendering a bird’s eye view, the entire model needs to be loaded into VRAM, limiting their scalability. Additionally, they may suffer from severe aliasing artifacts.
2.0.5 Anti-aliasing and Partition at Scale Dimension.
@@
Anti-aliasing has been a fundamental problem in computer graphics and image processing for a long time and has been extensively explored in the rendering community. The common approaches to tackle this problem can be categorized into two ways: super-sampling and pre-filtering. In the context of NeRF, super-sampling is casting multiple rays per pixel as in MobileNeRF [11], or sampling multiple points with perturbation, as demonstrated in Zip-NeRF [4]. This straightforward strategy is useful but also computationally expensive. On the other hand, Mip-NeRF [2] introduces the concept of pre-filtering through the integrated position encoding. Other research has explored scene partitioning at different scales to mitigate aliasing. For instance, BungeeNeRF [44] proposes a multi-scale MLP that progressively adds detail when the camera zooms in. However, akin to Mip-NeRF, it may only accurately represent the center of the scene, lacking detail elsewhere. Tri-MipRF [22] and D.Hu et al. [21] adopt a down-sampling strategy to create a smoother feature grid at each scale after training. PyNeRF [37] trained a multi-resolution hashed feature grid. These methods demonstrate impressive anti-aliasing capabilities and consistently high-quality rendering across the entire scene. However, their lack of space partitioning leads to challenges when scaling up, requiring the whole finest layer to render a single frame.
@@
3 Method
We present InfNeRF, a framework for efficient large-scale scene reconstruction using a LoD octree structure (Section 3.1). This tree structure naturally enables anti-aliasing rendering (Section 3.2), which significantly enhances the visual quality of rendered images. We also introduce tree pruning to improve model sparsity and compactness, along with a rendering approach tailored for the pruned structure (Section 3.3). Moreover, we propose an efficient training strategy for InfNeRF in Section 3.4. Finally, we also provide a comprehensive analysis of the computational complexity of InfNeRF in supplementary materials, offering insights into the efficiency and scalability of our approach in handling large-scale scene reconstructions.
3.1 Tree Structure
The core of our InfNeRF is a LoD structure, essentially an octree where each node represents a specific cubic space of the scene, known as the node’s Axis-Aligned Bounding Box (AABB). The octree is illustrated in the middle of Fig. 2. InfNeRF begins with the entire scene encapsulated within a root node, which is then divided into eight smaller cubes, corresponding to the child nodes. This division process continues recursively, partitioning the scene into increasingly finer parts, both in terms of space (size of the area) and scale (level of detail).
Each node in this tree is paired with its own NeRF, which represents the node’s AABB at a certain scale. Our approach is flexible regarding the choice of NeRF model. It can be a simple MLP [25], Instant-NGP [27], tri-plane [9, 22], or even a combination of them, as long as it can query the density and the color for a given 3D point and viewing direction :
@@ 3.1 Tree Structure
whose unit is in meters. The GSD, in this context, serves as a measure of resolution or granularity, indicating the level of detail that the node can capture about the physical world - the smaller the GSD, the finer the details. For example, if a node represents a 1 km cube with a grid size of 2048, the GSD will be meters. As each node in our octree shares the same grid size, the child nodes of have a GSD of meters, thereby capturing more detailed information. This is also illustrated in Fig. 2, where the GSD of a cube at level 1 is half that of a cube at level 0.
3.2 Anti-Aliasing Rendering
Similar to the original NeRF, the rendering process of InfNeRF accumulates the density and color of sampling points along a ray to determine the RGB of a pixel. However, the key difference is that the sampling points are queried from different nodes of our octree rather than a single NeRF as shown in Fig. 2. In the following, we show how we find the corresponding node for a given sampling point.
According to the sampling theory [41], a sample on the ray is not merely an infinitely small point, but rather a Gaussian sphere with a certain volume. For simplicity, we model this as a sphere with radius that corresponds to the pixel size,
as illustrated on the left of Fig. 2. Given a sampling sphere with position and radius , we aim to find the node whose granularity (GSD) matches and whose AABB contains . In this way, we ensure InfNeRF samples a proper region in the 3D space rather than just a single point, leading to a desirable anti-aliasing effect.
Since the GSD is discrete in our tree structure, we can either interpolate between two tree levels, as done in PyNeRF [37], or simply sample from the nearest level. InfNeRF opts for the latter, rounding towards the root for computational efficiency. Specifically, the level to sample a given sphere is determined by:

### v2

# InfNeRF: Towards Infinite Scale NeRF Rendering with Space Complexity

@@ Abstract.
In this work, we extend this proven LoD technique to Neural Radiance Fields (NeRF) by introducing an octree structure to represent the scenes in different scales. This innovative approach provides a mathematically simple and elegant representation with a rendering space complexity of , aligned with the efficiency of mesh-based LoD techniques. We also present a novel training strategy that maintains a complexity of . This strategy allows for parallel training with minimal overhead, ensuring the scalability and efficiency of our proposed method. Our contribution is not only in extending the capabilities of existing techniques but also in establishing a foundation for scalable and efficient large-scale scene representation using NeRF and octree structures. Code and checkpoints are available at: https://jiabinliang.github.io/InfNeRF.io/
figure description
1. Introduction
Recent progress in Neural Radiance Fields (NeRF) has significantly advanced the field of photo-realistic novel-view synthesisĀ (Mildenhall etĀ al., 2021; Barron etĀ al., 2021, 2023). Surpassing traditional 3D representations like meshes, NeRF demonstrates a superior capability in generating accurate novel views, particularly for transparent or reflective geometries. However, despite these advancements, a predominant focus of NeRF research has been confined to single objects or limited-scale scenes. The challenge of effectively representing extensive, large-scale scenes remains largely underexplored and unresolved.
@@ 1. Introduction
1.1. Block-NeRF and Mega-NeRF
@@
Towards this direction, recent large-scale NeRF approaches, such as Block-NeRFĀ (Tancik etĀ al., 2022) and Mega-NeRFĀ (Turki etĀ al., 2022) divide the whole neural field equally in space into smaller, more manageable blocks, referred to as grid partition in this paper. This allows the rendering device to load only the data relevant to the cameraās current location. However, these methods encounter challenges in scenarios requiring a birdās eye view of the entire scene. Such a panoramic view plays a crucial role in human interactive navigation. For example, to adjust the position of the camera, it is much more efficient to first zoom out, identify the target location, and then zoom in again, compared to the laborious process of navigating at a constant zoomed-in level. In such scenarios, the existing methods need to load all blocks simultaneously, which can overwhelm the memory capacity of the device. In addition to the memory issue, these approaches are prone to aliasing artifacts that arise from an insufficient sampling rate for high-frequency signals as illustrated in Fig. 1.
1.2. Proposed InfNeRF
Example of multi-resolution octree
Illustration of LoD in InfNeRF
@@ 1.2. Proposed InfNeRF
In this paper, we propose InfNeRF, a novel method that addresses the above challenges. The core concept of InfNeRF is inspired by the classical mesh-based Level of Detail (LoD) techniquesĀ (Clark, 1976; Erikson etĀ al., 2001; Ribelles etĀ al., 2010; Tanner etĀ al., 1998). An overview is illustrated in Fig. 2. Specifically, we construct a LoD tree for the target scene, where each node encapsulates a small NeRF with a roughly equivalent number of parameters. The root node provides a coarse representation of the entire scene, while its successive children progressively describe more detailed sub-regions. During rendering, each sampling point is forwarded by a different NeRF, depending on its location and radius. The density and color then are accumulated to determine the pixel color.
@@
As illustrated in Fig. 3, InfNeRF only requires a minimum subset of the nodes in rendering, chosen based on their proximity to the camera and their intersection within the frustum. This significantly reduces the memory burden and the I/O time for retrieving parameters from disk or cloud storage. The memory complexity is notably efficient at , where denotes the total information captured in the reconstruction. In the real world, with , we can represent the Earth down to a resolution of centimeters. In general, InfNeRF is like the combination of Mega-NeRFĀ (Turki etĀ al., 2022) and PyNeRF. It divides the scene in both spatial and scale dimensions while existing approaches typically only consider one of them.
@@
Another remarkable feature of InfNeRF is that the parent nodes in the LoD tree act as a smoother, low-pass-filtered version of the scene. Therefore, the proposed method naturally removes high-frequency components in distant views, which effectively reduces the aliasing artifacts and enhances the visual quality of the generated images (Fig.Ā 1). In our experiments, InfNeRF achieves superior rendering quality for large-scene reconstruction, with an improvement of over 2.4 dB in PSNR while accessing only 17% of the total parameters.
@@
In addition, the training of InfNeRF only increases slightly and maintains a time and space complexity of . Moreover, this strategy allows for the training process to be effectively parallelized across machines.
InfNeRF imposes no constraints on the underlying NeRF model and makes no assumptions about the scene, making it an adaptable framework that can seamlessly accommodate and leverage the rapid advancements in NeRF models.
@@
2. Related Work
2.1. Photogrammetry
@@
Among the established reconstruction techniques, mesh-based reconstruction methods called photogrammetry have been used widely for it efficiency and robustness in creating digital twins world. Open-source tools, such as COLMAPĀ (Schonberger and Frahm, 2016) and OpenMVSĀ (Cernea, 2020), have seamlessly integrated this workflow and gained widespread adoption in the academic community. Furthermore, commercial software solutions like ContextCapture and DJI TerraĀ (DJI, 2024) have significantly elevated the robustness and efficiency to meet industrial standards.
2.2. Level of Detail
@@
When the scene is large, it becomes impractical to store it entirely in memory. In this case, the Level of Detail (LoD) or Hierarchical Level of Detail (HLoD)Ā (Clark, 1976; Erikson etĀ al., 2001; Hoppe, 1998, 2023; Ulrich, 2002; Luebke, 2003) provides a framework for efficiently storing, fetching, and rendering the scene.
@@
Take tiled 2D mapĀ (Tiled web map, 2024) used widely in 2D navigation system for an example. The maps in different resolutions are cut into 256256 PNG images referred to as tiles. These tiles are organized into a quadtree structure, typically comprising 22 levels. This hierarchical arrangement allows web browsers to rapidly fetch data and render any location at any zoom level with constant space complexity.
@@
A similar technique can also be applied to 3D mesh. After the reconstruction of the scene into one detailed mesh, this mesh undergoes multiple simplifications Ā (Daniels etĀ al., 2008) to reduce the number of triangles. Concurrently, its texture also undergoes downsampling multiple timesĀ (Williams, 1983). Together this results in a mesh pyramid with varying levels of detail, ranging from fine to coarse. Subsequently, these meshes are partitioned into numerous tiles and organized hierarchically, resembling a tree structure similar to a quadtree in 2D mapping. During rendering, 3D engine such as CesiumJSĀ (CesiumJS, 2024) or OpenSceneGraphĀ (Burns and Osfield, 2004) walks the tree, then progressively fetch tiles based on their proximity and intersection with the frustum. This LoD technique enables users to have a seamless 3D navigation experience in a very large scene while maintaining a low memory footprint.
2.3. NeRF
@@
NeRF (Mildenhall et al., 2021) employs Multi-Layer Perceptrons (MLPs) for NeRF to capture a continuous density and viewpoint-dependent radiance of a scene. Methods like TensoRF (Chen et al., 2022) and MERF (Reiser et al., 2023) decompose the dense grid into low dimension feature tensor, while InstantNGP (Müller et al., 2022) opts to store these vectors in a sparse hash table. These strategies greatly enhance model compactness while maintaining speed.
2.4. Large-Scale Scene and Partition at Space Dimension
@@
While previous NeRF approaches primarily focus on scenes of limited scale, scaling up NeRF to handle large-scale scenes, such as cities or even the entire Earth, would empower a broader range of applications. Mip-NeRF 360Ā (Barron etĀ al., 2022) proposed a scene contraction technique to compress the unbounded scene into a bounded region. However, this approach compromises quality in the compressed areas. Grid-Guided NeRFĀ (Xu etĀ al., 2023) extends its capacity by combining a 2D feature grid and MLP, but its reliance on a 2D plane assumption limits its capability to handling complicated 3D scene. Block-NeRFĀ (Tancik etĀ al., 2022), Mega-NeRFĀ (Turki etĀ al., 2022), SMERFĀ (Duckworth etĀ al., 2023) and ScaNERFĀ (Wu etĀ al., 2023) divide the scene into several sub-regions and reconstruct them separately. In its specialized indoor scenario, SMERF can even achieve impressive real-time rendering. However, all these methods suffer from a common issue: when rendering a birdās eye view, the entire model has to be loaded into VRAM, limiting their scalability. Additionally, they may suffer from severe aliasing artifacts.
2.5. Anti-aliasing and Partition at Scale Dimension
@@
Anti-aliasing has been a fundamental problem in computer graphics and image processing for a long time and has been extensively explored in the rendering community. The common approaches to tackle this problem can be categorized into two ways: super-sampling and pre-filtering. In the context of NeRF, super-sampling is casting multiple rays per pixel as in MobileNeRFĀ (Chen etĀ al., 2023), or sampling multiple points with perturbation, as demonstrated in Zip-NeRFĀ (Barron etĀ al., 2023). This straightforward strategy is useful but also computationally expensive. On the other hand, Mip-NeRF Ā (Barron etĀ al., 2021) introduces the concept of pre-filtering through the integrated position encoding. Other research has explored scene partitioning at different scales to mitigate aliasing. For instance, BungeeNeRFĀ (Xiangli etĀ al., 2022) proposes a multi-scale MLP that progressively adds detail when the camera zooms in. Tri-MipRFĀ (Hu etĀ al., 2023a), PyNeRFĀ (Turki etĀ al., 2023) and Mip-VoGĀ (Hu etĀ al., 2023b) train a multi-resolution feature grid. These methods demonstrate impressive anti-aliasing capabilities and consistently high-quality rendering across the entire scene. However, their lack of space partitioning leads to challenges when scaling up, requiring the whole finest layer to render a close-up frame.
@@
2.6. Scene Partition with an Octree
Some researches, like PlenOctreesĀ (Yu etĀ al., 2021) and NGLoDĀ (Takikawa etĀ al., 2021) utilize an octree-based approach to represent the scene. PlenOctrees stores the spherical harmonics coefficients in the leaves and uses the octree to accelerate accessing, while in NGLoD, a signed distance field is stored in the octree to represent a LoD closed volume. These direct storage techniques provide fast access and training but require significant memory, limiting their scalability. In contrast, InfNeRF employs many compact hash tables as the underlying data storage. The shallow octree is primarily used for organizing the hash tables into a LoD structure. This approach results in a much more compact LoD representation and much better scalability.
2.7. Gaussian Splatting
In recent times, Gaussian splattingĀ (Kerbl etĀ al., 2023) has gained significant attention due to its fast rendering speed compared to NeRF. Additionally, the introduction of Mip-SplattingĀ (Yu etĀ al., 2024) effectively reduces aliasing artifacts across different resolutions. Then Hierarchical GaussiansĀ (Kerbl etĀ al., 2024), VastGaussianĀ (Lin etĀ al., 2024), and CityGaussianĀ (Liu etĀ al., 2024), explore its potential in large scene reconstruction. While Gaussian splatting may not achieve the same PSNR as traditional NeRFĀ (Liu etĀ al., 2024) in challenging scenarios such as varying light conditions or camera exposure, it remains a promising technique for large-scale scene reconstruction in the future.
3. Method
We present InfNeRF, a framework for efficient large-scale scene reconstruction using a LoD octree structure (SectionĀ 3.1). This tree structure naturally enables anti-aliasing rendering (SectionĀ 3.2), which significantly enhances the visual quality of rendered images. We also introduce tree pruning to improve model sparsity and compactness, along with a rendering approach tailored for the pruned structure (SectionĀ 3.3). Moreover, we propose an efficient training strategy for InfNeRF in SectionĀ 3.4. Finally, we provide a comprehensive analysis of the computational complexity of InfNeRF in supplementary material, offering insights into the efficiency and scalability of our approach in handling large-scale scene reconstructions.
3.1. Tree Structure
The core of our InfNeRF is a LoD structure, essentially an octree where each node represents a specific cubic space of the scene, known as the nodeās Axis-Aligned Bounding Box (AABB). The octree is illustrated in the middle of Fig. 2. InfNeRF begins with the entire scene encapsulated within a root node, which is then divided into eight smaller cubes, corresponding to the child nodes. This division process continues recursively, partitioning the scene into increasingly finer parts, both in terms of space (size of the area) and scale (level of detail).
Each node in this tree is paired with its own NeRF, which represents the nodeās AABB at a certain scale. Our approach is flexible regarding the choice of NeRF model. It can be a simple MLPĀ (Mildenhall etĀ al., 2021), Instant-NGPĀ (Müller etĀ al., 2022), triplaneĀ (Chan etĀ al., 2022; Hu etĀ al., 2023a), or even a combination of them, as long as it can query the density and the color for a given 3D point and viewing direction :
@@ 3.1. Tree Structure
whose unit is in meters. The GSD, in this context, serves as a measure of resolution or granularity, indicating the level of detail that the node can capture about the physical world - the smaller the GSD, the finer the details. For example, if a node represents a 1 km3 cube with a grid size of 2048, the GSD will be meters. As each node in our octree shares the same grid size, the child nodes of have a GSD of meters, thereby capturing more detailed information. This is also illustrated in Fig.Ā 2, where the GSD of a cube at level 1 is half that of a cube at level 0.
3.2. Anti-Aliasing Rendering
Similar to the original NeRF, the rendering process of InfNeRF accumulates the density and color of sampling points along a ray to determine the RGB of a pixel. However, the key difference is that the sampling points are queried from different nodes of our octree as shown in Fig.Ā 2 rather than a single NeRF . In the following, we show how we find the corresponding node for a given sampling point.
According to the sampling theoryĀ (Williams, 1983), a sample on the ray is not merely an infinitely small point, but rather a Gaussian sphere with a certain volume. For simplicity, we model this as a sphere with radius that corresponds to the pixel size,
as illustrated on the left of Fig. 2. Then, given a sampling sphere with position and radius , we aim to find the node whose granularity (GSD) matches and whose AABB contains . In this way, we ensure InfNeRF samples a proper region in the 3D space rather than just a single point, leading to a desirable anti-aliasing effect.
We observe that when the sampling sphere falls between two discrete levels, small aliasing occurs. Nevertheless when sampling points transition between levels, discrepancies in density or color may arise, engendering an undesirable āpopupā effect when camera moves. To mitigate these phenomenons, both PyNeRF and Tri-MipRF employ interpolation techniques between adjacent levels which will double the computation. However, we pursue a simpler approach by introducing stochastic perturbations to the radius of the sampling sphere.
This strategy use human eyeās natural blending effect to attenuate the abrupt transitions without incurring additional computational cost. Finally, the level to sample a given sphere is determined by: